OpenStudy (anonymous):

Find the average rate of change of the function over the given interval. Compare the average rate of change with the instant rate of change at the endpoints of the interval. f(x)=-1/x [1,2]

OpenStudy (anonymous):

f(x)=-1x^-1 -1(-1)x^-2 1/x^2

OpenStudy (anonymous):

1/(1)^2=1 1/(2)^2=2 instantaneous rates: f'(1)=1 ; f'(2)= 1/4

OpenStudy (anonymous):

how do i find the average rate?

OpenStudy (anonymous):

i know the average rate is 1/2 because the back of the book tells me, but I do not see how to get that from my work. :(

OpenStudy (dumbcow):

think of slope between the 2 endpoints avg rate = f(2) -f(1) / 2-1 = -1/2 -(-1) / 1 = 1/2

OpenStudy (anonymous):

so basically f(1)-f(2)?

OpenStudy (dumbcow):

or more technically using calculus \[avg rate = \frac{1}{b-a}\int\limits_{a}^{b}f'(x) dx = \frac{f(b) -f(a)}{b-a}\]

OpenStudy (anonymous):

im trying to understand really

OpenStudy (anonymous):

so if i take the original equation f(x)=-1/x

OpenStudy (anonymous):

then i do the f(2)-f(1)

OpenStudy (anonymous):

wait +f1)

OpenStudy (anonymous):

im just confusing myself now lol

OpenStudy (dumbcow):

yeah f(2) -f(1)

OpenStudy (anonymous):

-1/2-(-1/1)=1/2

OpenStudy (anonymous):

thanks

OpenStudy (dumbcow):

welcome